OpenStudy (anonymous):

Toby has written 65 poems. His brother has written more than 25 poems. How many poems have they written altogether? A. x + 25 ≤ 65 B. x – 25 > 65 C. 25 + x = 65

OpenStudy (study312):

I think it is A because \[x + 25 \le 65\] means x + 25 is more than or equal to 65 and if Toby wrote 65 and his brother wrote more than 25 then combined they would make a lot. This is a guess

OpenStudy (anonymous):

Thanks Can You Help With Another One?

OpenStudy (study312):

sure

OpenStudy (anonymous):

Nancy has a budget of $65. A ticket to a theme park costs at least $25. How much money will she have left after going to the theme park? A. x + 25 ≤ 65 B. x – 25 > 65 C. 25 + x = 65

OpenStudy (study312):

B because she is losing $25

OpenStudy (anonymous):

Can You Help Me With Two More Questions?

OpenStudy (study312):

how wait no! Its A because \[x - 25 \le 65\] means x - 25 is less or equal to 65. Sorry that means the first question is B

OpenStudy (study312):

I always get them mixed up, my teacher told me pretend that the symbol is a fish mouth going to eat the bigger number

OpenStudy (anonymous):

Ok Thanks

OpenStudy (study312):

Whats the next question?

OpenStudy (anonymous):

Arvee biked a total of 25 miles in September. He biked more miles in October. If he rode 65 miles in all, how many miles did he ride in October? A. x + 25 ≤ 65 B. x – 25 > 65 C. 25 + x = 65

OpenStudy (study312):

C because he biked 65 miles in all and x is the miles he biked in October. I know this is right

OpenStudy (study312):

Is that all?

OpenStudy (anonymous):

Ok Last Question Francis planted 65 flowers. She planted 25 lilies and some tulips. How many tulips did she plant? A. x + 25 ≤ 65 B. x – 25 > 65 C. 25 + x = 65

OpenStudy (study312):

well its not asking for it but 65 - 25 = 40 there is 40 tulips. C because she planted 65 in all and x is the number of tulips

OpenStudy (anonymous):

Ok Thank You. That Was It ^-^

OpenStudy (study312):

Man this is really repetitive, they ask the same question 4 times. I guess they really want you to get it lol

OpenStudy (anonymous):

I Know Right xDD Anyways Thanks ^-^